OpenStudy (olivia567):

In an L-C-R circuit, capacitance is changed from C to 2C. For the resonant frequency to remain unchanged, the inductance should be changed form L to ________ ? A) 4L B)2L C)L/2 D)L/4

OpenStudy (olivia567):

Answer is L/2.. But why is it so?? Can anyone explain?

OpenStudy (irishboy123):

\[w_o = 2 \pi f = \frac{1}{\sqrt{LC}}= \frac{1}{\sqrt{(L^*/2)(2C^*)}}\]

OpenStudy (shaswatrajput):

\[wo=1/\sqrt(L*C)\] so, if C changes from C to 2C, inductance will change from L to L/2